OpenStudy (anonymous):

A car drives down a road in such a way that its velocity ( in m/s) at time t (seconds) is v(t) = 1 t^{1/2} + 3 Find the car's average velocity (in m/s) between t=3 and t=6.

OpenStudy (anonymous):

i think you need the total distance travelled divided by the time elapsed

OpenStudy (anonymous):

i.e.i think you need \[\frac{1}{3}\int_3^6 \sqrt{t}+3 dt\]

OpenStudy (anonymous):

\[\frac{1}{3}\int_3^6 \sqrt{t}dt+\frac{1}{3}\times 9\] \[\frac{1}{3}(4\sqrt{6}-2\sqrt{3})+3\]

OpenStudy (anonymous):

thank u ! , i have been on that question for an hr!